# LLaVA-PruMerge: Adaptive Token Reduction for Efficient Large Multimodal Models

### v2

@@ 3.2 Adaptive Important Token Selection via Outlier Detection
Algorithm 1 Token PruMerge algorithm for reducing the number of visual tokens.
@@
6:for $p=\{i_{1},\cdots,i_{m}\}$ do
7:     Calculate the distance between selected token $\mathbf{y}_{p}$ and other visual tokens, $\mathbf{y}_{\{1,\cdots,n\}/p}$;
8:     Use $\mathbf{y}_{p}$ as cluster center and run $k$-nearest neighbor algorithm to find $k$ similar tokens, with indices $\{j_{1},\cdots,j_{k}\}_{p}$;
9:     Update cluster center token with weighted sum: $\mathbf{y}_{p}^{\prime}=\sum_{q=1}^{k}\mathbf{a}[j_{q}]\cdot\mathbf{y}_{j_{q}}$;
10:end for
11:(see Sec. 3.3).
12:Output a refined stack of visual token $\mathbf{Y}^{\prime}=\{\mathbf{y}_{1}^{\prime},\cdots\mathbf{y}_{m}^{\prime}\}$.
@@ 4 Experiments
LLaVA-1.5 + Ours Vicuna-7B 336 558K 665K 72.0 68.5 56.0 86.3 1350.3 60.9
@@
LLaVA-1.5 + Ours Vicuna-13B 336 558K 665K 72.8 71.0 58.4 86.2 1428.2 62.3
@@ 4.3.2 Effectiveness of Each Module in PruMerge
LLaVA-1.5 w/ AITS Vicuna-7B 66.5 54.8 85.7 1221.6
LLaVA-1.5 w/ AITS & TS Vicuna-7B 68.5 56.0 86.3 1350.3
@@
LLaVA-PruMerge w. fine-tuning Vicuna-7B 68.5 56.0 86.3 1350.3
@@ 5 Conclusion
6 Limitations and Future Work
Our exploration of LLaVA-PruMerge has two primary limitations. First, the visual token compression, while efficient, is not entirely lossless. This results in a marginal performance gap between the original model (LLaVA) and our optimized version LLaVA-PruMerge. We are committed to advancing our research towards achieving a fully lossless token compression algorithm, aiming to close these performance gaps. Second, the scope of our validation efforts is somewhat constrained by the computational resources typically available in academic settings. This limitation has precluded a comprehensive assessment of PruMerge’s applicability to larger-scale models, such as those envisioned in the LLaVA-Next [Liu et al., 2023a] framework with a Yi-34B backbone. Future investigations will seek to extend our methodology to these more expansive model architectures, exploring its potential for generalization and broader impact.

### v3

@@ 3.2 Adaptive Important Token Selection via Outlier Detection
Algorithm 1 Token PruMerge and PruMerge+ algorithms for reducing the number of visual tokens.
@@
6:(Optional for PruMerge+) Calculate the outlier ratio $r_{o}=\frac{m}{n}$.
7:(Optional for PruMerge+) Spacial-uniformly sample visual tokens with $r_{o}=\frac{m}{n}$, and get $\{i_{1+m},\cdots,i_{2m}\}$.
8:(Optional for PruMerge+) Update the selected tokens index with $\{i_{1},\cdots,i_{m},1_{1+1},\cdots,i_{2m}\}$.
9:for $p=\{i_{1},\cdots,i_{m}\}$ do
10:     Calculate the distance between selected token $\mathbf{y}_{p}$ and other visual tokens, $\mathbf{y}_{\{1,\cdots,n\}/p}$;
11:     Use $\mathbf{y}_{p}$ as cluster center and run $k$-nearest neighbor algorithm to find $k$ similar tokens, with indices $\{j_{1},\cdots,j_{k}\}_{p}$;
12:     Update cluster center token with weighted sum: $\mathbf{y}_{p}^{\prime}=\sum_{q=1}^{k}\mathbf{a}[j_{q}]\cdot\mathbf{y}_{j_{q}}$;
13:end for
14:(see Sec. 3.3).
15:Output a refined stack of visual token $\mathbf{Y}^{\prime}=\{\mathbf{y}_{1}^{\prime},\cdots\mathbf{y}_{m}^{\prime}\}$.
@@ 3.3 Token Supplement via Similar Key Clustering
3.4 PruMerge+: Bridging the Efficiency-Performance Gap
While PruMergeachieves a remarkable reduction in the number of visual tokens—over tenfold compared to the original setup—the process is not without its drawbacks. Specifically, the compression technique, though efficient, introduces a marginal performance discrepancy between the LLaVA model and its PruMerge-optimized counterpart, LLaVA-PruMerge. To address this, we introduce PruMerge+, a refined version that strikes an optimal balance between token reduction and model performance.
Figure 4: Image Visualizations of PruMerge and PruMerge+.
PruMerge+ enhances our original method by maintaining the ability to significantly reduce visual token count—by an average of fourfold—without any resultant performance degradation. This improvement is detailed in Algorithm 1, building upon the token selection strategies outlined in Section 3.2. A new aspect of PruMerge+ lies in its enhanced token selection process. Beyond merely focusing on the previously identified important tokens, PruMerge+ extends its reach to encompass additional visual tokens from areas initially deemed less critical. This is achieved through a spatially uniform sampling of visual tokens, guided by a predetermined ratio informed by the distribution of outlier tokens. This methodology ensures a more comprehensive and representative selection of visual tokens, as depicted in Figure 4, thereby minimizing any performance losses while still achieving substantial token reduction.
@@ 4 Experiments
LLaVA-1.5 + PruMerge Vicuna-7B 336 558K 665K 72.0 68.5 56.0 76.3 1350.3 60.9
LLaVA-1.5 + PruMerge+ Vicuna-7B 336 558K 665K 76.5 68.3 57.1 84.0 1431.6 63.5
@@
LLaVA-1.5 + PruMerge Vicuna-13B 336 558K 665K 72.8 71.0 58.4 78.5 1428.2 62.3
LLaVA-1.5 + PruMerge+ Vicuna-13B 336 558K 665K 77.8 71.0 58.6 84.4 1485.5 65.7
@@ 4.3.2 Effectiveness of Each Module in PruMerge
LLaVA-1.5 w/ AITS Vicuna-7B 66.5 54.8 75.7 1221.6
LLaVA-1.5 w/ AITS & TS Vicuna-7B 68.5 56.0 76.3 1350.3
@@
LLaVA-PruMerge w. fine-tuning Vicuna-7B 68.5 56.0 76.3 1350.3